# Fourfold path to Thermality: Inequivalent purifications of Rindler wedge

### v1

@@ II.1.4 Path 4: Null Shifts via Rindler Wedge $R_{4}$
Path 4 involves the wedge $R_{4}$, whose origin is obtained from $R_{1}$ by a null shift along the $U_{1}$-axis by parameter $\Delta_{4}$. Subsequently, $R_{3}$ is reached by an additional null shift from $R_{4}$ along the $V_{2}$-axis with parameter $\Delta_{3}$, as depicted in Figure 2. The wedge $R_{4}$ is an independent Rindler wedge analogous to $R_{2}$, obtained by shifting through a different null axis, and plays a central role in the construction of this path.
These construction paths serve to systematically characterise inequivalent purifications of Rindler vacua and reveal how selective and full thermalisation arise in distinct regions of spacetime. While intermediate wedges are omitted in some scenarios for simplicity, the analysis extends to more general initial states, though with potentially differing particle spectra [8].
@@ II.2 Coordinate Relations among the key players
Two dimensional Minkowski spacetime is denoted as $\mathcal{M}$, parametrized by inertial coordinates $(X,T)$, and a hierarchy of Rindler wedges  [19], denoted $(R_{i})$, each described by coordinates $(x_{i},t_{i})$ for $i\in\{1,2,3,4\}$. As depicted in Figs . 1 and  2, the nested structure of Rindler wedges are understood as inclusions of spacetime regions determined by their null horizons, such that $R_{3}\subset R_{2}\subset R_{1}\subset M$, and $R_{3}\subset R_{4}\subset R_{1}\subset M$, with individual wedges distinguished by the shaded regions in pale red, blue, and green. Wedges $R_{2}$ and $R_{3}$ are generated from $R_{1}$ by successive null displacements along the relevant Minkowski light rays, a construction central to our analysis.
II.2.1 Various Rindler wedges and their coordinate relationships
@@
II.2.2 Minkowski to Rindler-3 and Rindler 1 to Rindler-3 via spatial shift along $x_{1}$ axis
@@
where $\Delta_{1}$ encodes the coordinate shift in the Rindler space, along $x_{1}$ axis relates the wedge $R_{3}$ to $R_{1}$. And the lightcone coordinates in Minkowski spacetime are related to the coordinates of $R_{3}$ as:
@@
II.2.3 Rindler 1 to Rindler 3 via Rindler 2: series of null shifts
@@
We now relate Rindler-3 ($R_{3}$) to Minkowski space. Its transformation is given by:
@@
II.2.4 Rindler 1 to Rindler 3 via Rindler 4: series of null shifts
@@
These expressions completely characterise the relationships among the Rindler patches that shift to the null in the $U$ and $V$ directions.
These construction paths serve to systematically characterise inequivalent purifications of Rindler vacua and reveal how selective and full thermalisation arise in distinct regions of spacetime. While intermediate wedges are omitted in some scenarios for simplicity, the analysis extends to more general initial states, though with potentially differing particle spectra [8].
Having established the coordinate relations among the Rindler wedges and the four inequivalent paths leading to $R_{3}$, we now turn to the operational question: what does an observer in $R_{3}$ actually measure along each path? In particular, we diagnose thermality using two complementary criteria,(i) the expectation value of particle number operators defined with respect to $R_{3}$ modes, and (ii) the expectation value of the stress tensor computed via the Virasoro anomaly. These diagnostics allow us to distinguish Gibbsian thermality arising from entanglement from non-Gibbsian thermality arising purely from Bogoliubov mixing.
@@ III Methodology
In the previous section, we defined the coordinate systems in various spacetimes relevant to the problem and then obtained the relations between them. To determine the particle content in various spacetimes, we use two techniques. One is the canonical way to set up the mode expansion of scalar fields and evaluate the Bogoliubov coefficient. Another useful method is using the Virasoro anomaly.
III.1 Bogolibov method
@@
The normalisation constant $N_{\omega}$ is determined by requiring orthonormality of the modes under the Klein–Gordon inner product, defined on a constant-time hypersurface $\Sigma$ as [7]
@@
for $l\neq m=1,2$. From this normalisation condition, one obtains
@@
We now introduce the mode expansions of the scalar field in the various Rindler wedges relevant to our analysis. These modes represent the natural positive-frequency excitations with respect to the proper time of uniformly accelerated observers associated with each wedge. Since thermality in Rindler spacetime is defined operationally through the response of such observers, particle content must be evaluated by comparing these mode decompositions across different wedges. Although several Rindler frames are introduced below, in each path, only a specific subset of modes undergoes nontrivial Bogoliubov mixing. Crucially, in two dimensions the field separates into independent right- and left-moving sectors, allowing us to track how thermal behaviour may arise in one chiral sector while the other remains in its vacuum. This chiral structure will play a central role in distinguishing full Gibbsian thermality from the selective, non-Gibbsian thermality induced by null shifts. The right-moving component in wedge $R_{1}$ takes the form
@@
while in the remaining wedges the expansions are
@@
III.2 Bogoliubov Transformation and the Number Operator
@@
III.3 Virasoro anamoly method
In two dimensions, all components of $\langle T_{\mu\nu}\rangle$ can be determined using the Virasoro anomaly of the normal-ordered stress tensor. While normal ordering is well defined in Minkowski spacetime, its extension to curved spacetime can be motivated by applying it in locally inertial frames, which exist at each spacetime point. This prescription is consistent with general covariance and enables the determination of all components of the covariant quantum stress tensor.
A potential subtlety arises from the use of local coordinates, since the definition of a quantum state is intrinsically global. In particular, normal ordering with respect to positive-frequency modes $\frac{1}{\sqrt{4\pi\omega}}e^{-i\omega u}$ and $\frac{1}{\sqrt{4\pi\omega}}e^{-i\omega v}$ defined in local coordinates may appear ambiguous from a global standpoint. In semiclassical gravity, this issue is addressed by restricting to Hadamard states, for which the short-distance singularity structure of the two-point function is locally Minkowskian and admits a consistent renormalisation prescription [28, 3, 22].
In this work, we employ Rindler coordinates, which cover only a portion of Minkowski spacetime. Nonetheless, the normal-ordered stress tensor is well defined within the Rindler wedge and corresponds to renormalisation with respect to the Rindler vacuum[6]. The resulting expectation values have a clear operational interpretation in terms of observables accessible to uniformly accelerated observers and are directly related to the Unruh effect [24, 25, 26]. This restriction to the Rindler wedge does not compromise the consistency of the semiclassical description.
Under conformal transformation $f(u)$, the energy-momentum tensor behaves as
$T^{\prime}(u)=\bigg(\frac{\partial f}{\partial u}\bigg)^{2}T\big(f(u)\big)+\frac{c}{12}S(U,u)$
Where $S(U,u)$ denotes the Schwarzian derivative and is defined as[4]:
@@ III.3 Virasoro anamoly method
And, $c$ denotes the central charge.
@@ IV Particle content in $R_{3}$ along various paths
In this section, we analyse, in a unified and comparative manner, the particle content detected by an accelerated observer confined to the fixed Rindler wedge $R_{3}$ along four inequivalent construction paths. In each case, the notion of particles is defined with respect to the positive-frequency Rindler modes adapted to $R_{3}$, and thermality is diagnosed operationally through the expectation values of the corresponding number operators. While all four paths can give rise to thermal particle spectra in $R_{3}$, the physical origin of this thermality differs markedly among them. In particular, we distinguish between thermality arising from horizon-induced entanglement, which leads to a mixed Gibbs state and thermalisation of both chiral sectors, and thermality generated purely by Bogoliubov mixing under null-shift transformations, which produces a thermal spectrum in only one chiral sector while the global and local quantum states remain pure. By keeping the final wedge $R_{3}$ fixed and varying only the global state and the sequence of coordinate transformations, we isolate the mechanisms responsible for full versus selective thermalisation in Rindler spacetime.
@@ IV.1 Particle content along path 1
In this section, we evaluate the particle content along Path 1, focusing on the particle content in $R_{3}$ when the field is in the Minkowski vacuum. The wedge $R_{3}$ is spatially shifted from the origin by a parameter $\Delta_{1}$, as illustrated in Figs. 1 and 2. Since the Unruh effect is invariant under such spatial translations, we expect the particle content in $R_{3}$ to be thermal. We demonstrate this explicitly below.
@@
Thus, the right-moving modes in wedge $R_{3}$, when evaluated in the Minkowski vacuum, are thermally excited and exhibit a Planck distribution. A similar calculation shows that the left-moving modes also thermalize, confirming full thermality along Path 1.
@@
For subsequent sections, we take the initial state to be the vacuum $|0_{R_{1}}\rangle$ defined with respect to positive frequency Rindler modes localized entirely in the right wedge $R_{1}$:
$\overrightarrow{\hat{a}}_{1}(\omega)|0_{R_{1}}\rangle=0,\quad\overleftarrow{\hat{b}}_{1}(\omega)|0_{R_{1}}\rangle=0.$ (80)
Here, $\overrightarrow{\hat{a}}_{1}(\omega)$ and $\overleftarrow{\hat{b}}_{1}(\omega)$ annihilate right- and left-moving modes in $R_{1}$, respectively.
For an $R_{1}$ observer, the expectation values of particle number operators vanish:
$\langle 0_{R_{1}}|\overrightarrow{\hat{N}}_{R_{1}}(\omega)|0_{R_{1}}\rangle=0,\quad\langle 0_{R_{1}}|\overleftarrow{\hat{N}}_{R_{1}}(\omega)|0_{R_{1}}\rangle=0.$
Since the left wedge degrees of freedom are excluded in our construction, the $|0_{R_{1}}\rangle$ state remains pure and unentangled, with no partial trace performed. Hence, the standard Unruh mechanism involving tracing over inaccessible modes does not operate here [9]. We adopt this scenario as a baseline: in the $R_{1}$ vacuum, an observer detects no particles, and thermality observed later is defined relative to this baseline.
@@ IV.2 Particle content along path 2
In this section, we evaluate the particle content along Path 2, which connects the Rindler wedges $R_{1}$ and $R_{3}$ through a spatial shift parametrized by $\Delta_{1}$, as illustrated in Figs. 1 and 2. To compute the Bogoliubov coefficients relating the mode functions in these two wedges, we proceed as follows.
@@
To calculate the required Bogoliubov coefficients, we begin by multiplying both sides of Eq. (47) by the Fourier kernel $\int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}e^{i\Omega^{\prime}u_{2}}$. After simplifications and using the orthogonality of exponentials, we obtain [16]
$\begin{split}\int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}\;e^{i\;\Omega^{\prime}\;u_{2}}\;\hat{\phi}_{2}=\int_{0}^{\infty}\frac{d\Omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\Omega}}\\ \int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}\;\bigg(\overrightarrow{\hat{a}}_{2}(\Omega)\;e^{i\;(\Omega^{\prime}-\Omega)\;u_{2}}+\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\;e^{i\;(\Omega+\Omega^{\prime})\;u_{2}}\bigg),\end{split}$ (95)
@@
Due to orthogonality, valid for $\Omega^{\prime}>0$.
Next, we compute the same integral using the field operator $\hat{\phi}_{1}(u_{1})$, expressed in terms of $u_{2}$ via the null-shift relation connecting wedges $R_{1}$ and $R_{2}$. Multiplying both sides of Eq. (46) by the same Fourier kernel gives
$\begin{split}\int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}\;e^{i\;\Omega^{\prime}\;u_{2}}\;\hat{\phi}_{1}=\frac{1}{2\pi}\int_{0}^{\infty}\frac{d\omega}{\sqrt{2\omega}}\bigg(\overrightarrow{\hat{a}}_{1}(\omega)\int_{-\infty}^{\infty}du_{2}\\ \;e^{i\;\Omega^{\prime}\;u_{2}-i\;\omega\;u_{1}}+\overrightarrow{\hat{a}}_{1}^{\dagger}(\omega)\;\int_{-\infty}^{\infty}du_{2}\;e^{i\;\Omega^{\prime}\;u_{2}+i\;\omega\;u_{1}}\bigg),\end{split}$ (97)
@@
The Rindler vacuum in wedge $\mathrm{R}_{1}$ is defined by the annihilation condition
$\overleftarrow{\hat{b}}_{1}|0_{R_{1}}\rangle=0,$ and factorises as $|0_{R_{1}}\rangle=|\overleftarrow{0}_{R_{1}}\rangle\otimes|\overrightarrow{0}_{R_{1}}\rangle,$ where $|\overleftarrow{0}_{R_{1}}\rangle$ and $|\overrightarrow{0}_{R_{1}}\rangle$ denote the left- and right-moving vacuum states, respectively. The operators $\overleftarrow{\hat{b}}_{1}$ and $\overleftarrow{\hat{b}}_{1}^{\dagger}$ satisfy canonical commutation relations analogous to those in Eq. (59).
@@
This integral is analytically tractable (see Appendix B) and evaluates to
@@
[…]
@@
This confirms that the left-moving modes are thermally excited during the transition from $\mathrm{R}_{1}$ to $\mathrm{R}_{2}$, at a characteristic Unruh-like temperature $T=\frac{a}{2\pi}$. In contrast, the right-moving modes remain in the vacuum state.
@@
Since $\mathrm{R}_{2}$ is obtained from $\mathrm{R}_{1}$ through a null shift, the resulting quantum field configuration exhibits selective thermalisation, the left-moving modes become thermally populated, while the right-moving modes remain in vacuum, as previously established. The corresponding quantum state in $\mathrm{R}_{2}$ factorizes as a tensor product in the Hilbert space $\mathcal{H}_{R_{2}}$ and is expressed as [6]
$\mid\psi_{R_{2}}\rangle=\left(\sum_{n=0}^{\infty}\sqrt{p_{n}}\mid\overleftarrow{n}\rangle\right)\otimes\mid\overrightarrow{0}\rangle,$ (121)
where $p_{n}=\frac{e^{-n\beta\Omega}}{Z},\quad Z=\sum_{n=0}^{\infty}e^{-n\beta\Omega},$ denote the thermal weights and partition function, ensuring $\mid\psi_{R_{2}}\rangle$ is normalized [11, Ch. I, §1.3-1.4], [17, Ch. 7]. Here, $\mid\overleftarrow{n}\rangle$ is the $n$-particle Fock basis state of the left-moving mode, and $\mid\overrightarrow{0}\rangle$ is the vacuum state of the right-moving mode in $\mathrm{R}_{2}$. A detailed discussion of this state’s structure and its relation to thermal density matrices and purification is provided in Appendix C.
We now compute the expectation value of the number operator for right-moving modes in $\mathrm{R}_{3}$ evaluated in the state $\mid\psi_{R_{2}}\rangle$. The number operator
@@
$\langle\overrightarrow{\hat{N}}_{32}(\nu)\rangle=\langle\psi_{R_{2}}\mid\overrightarrow{\hat{a}}_{3}^{\dagger}(\nu)\overrightarrow{\hat{a}}_{3}(\nu)\mid\psi_{R_{2}}\rangle.$ (122)
@@
$\begin{split}\langle\overrightarrow{\hat{N}}_{32}(\nu)\rangle=\int_{0}^{\infty}d\Omega\int_{0}^{\infty}d\Omega^{\prime}\,\overrightarrow{\beta^{*}_{32}}(\nu,\Omega)\overrightarrow{\beta}_{32}(\nu,\Omega^{\prime})\\ \times\langle\psi_{R_{2}}\mid\overrightarrow{\hat{a}}_{2}(\Omega^{\prime})\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\mid\psi_{R_{2}}\rangle.\end{split}$ (123)
Since $\mid\psi_{R_{2}}\rangle$ is a tensor product with the right-moving mode in vacuum, the two-point function reduces to
$\begin{split}\langle\psi_{R_{2}}\mid\overrightarrow{\hat{a}}_{2}(\Omega^{\prime})\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\mid\psi_{R_{2}}\rangle=\langle\overrightarrow{0}\mid\overrightarrow{\hat{a}}_{2}(\Omega^{\prime})\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\mid\overrightarrow{0}\rangle\\ =\delta(\Omega-\Omega^{\prime}).\end{split}$ (124)
@@
The factor $\delta(0)$ arises from the mode normalization in the continuum and represents an infinite constant related to the volume. Hence, the mean particle number density per unit frequency $\nu$ in $\mathrm{R}_{3}$, as measured in $\mathrm{R}_{2}$, is
@@
expressing the familiar Planckian distribution at the Unruh like temperature $T=\frac{a}{2\pi}$.
We conclude that the right-moving modes mediating between $\mathrm{R}_{2}$ and $\mathrm{R}_{3}$ exhibit a thermal population characterized by this Planckian spectrum when the field in $\mathrm{R}_{2}$ is selectively excited state $\mid\psi_{R_{2}}\rangle$.

### v2

@@ II.1.4 Path 4: Null Shifts via Rindler Wedge $R_{4}$
Path 4 involves the wedge $R_{4}$, whose origin is obtained from $R_{1}$ by a null shift along the $U_{1}$-axis by parameter $\Delta_{4}$. Subsequently, $R_{3}$ is reached by an additional null shift from $R_{4}$ along the $V_{2}$-axis with parameter $\Delta_{3}$, as depicted in Figure 2. The wedge $R_{4}$ is an independent Rindler wedge analogous to $R_{2}$, obtained by shifting through a different null axis. The idea is the same as that of path 3 except that wedge $R_{2}$ is replaced with wedge $R_{4}$. Again, we note that the starting point of path 4 could be that the wedge $R_{4}$ has a right-moving thermal flux of radiation, whereas there is no particle content in the right-moving sector. For this initial configuration as well, the particle distribution in $R_{3}$ is thermal. Although this is possible, in this work, we focus on an analysis that begins with the Rindler wedge $R_{1}$ in the vacuum state.
@@ II.2 Coordinate Relations among the key players
In this subsection, we introduce the various coordinate systems relevant to our setup and describe the relations among them.
II.2.1 Minkowski spacetime to Rindler $1$ wedge
@@
II.2.2 Minkowski to Rindler-3 and Rindler-1 to Rindler-3 via spatial shift along $x_{1}$ axis
@@
where $\Delta_{1}$ encodes the coordinate shift in the Minkowski spacetime, along $x_{1}$ axis relates the wedge $R_{3}$ to $R_{1}$ ,if possible we can shown $\Delta_{1}$ in figure. The lightcone coordinates in Minkowski spacetime are related to the coordinates of $R_{3}$ as:
@@
II.2.3 Rindler-1 to Rindler-3 via Rindler-2: series of null shifts
@@
We now relate the Rindler-3 ($R_{3}$) coordinate system to Minkowski spacetime coordinates. Its transformations are given by:
@@
II.2.4 Rindler-1 to Rindler-3 via Rindler-4: series of null shifts
@@
These expressions completely characterize the relationships among the Rindler patches that shift to the null in the $U$ and $V$ directions.
These construction paths serve to systematically characterize inequivalent purifications of Rindler vacua and reveal how selective and full thermalization arise in distinct regions of spacetime. While intermediate wedges are omitted in some scenarios for simplicity, the analysis extends to more general initial states, though with potentially differing particle spectra [7].
@@ III Methodology
In the previous section, we defined the coordinate systems in various spacetimes of relevance and obtained the relations between them. To determine the particle content in various spacetimes, we use two techniques. One is the canonical way to set up the mode expansion of scalar fields and to evaluate the Bogoliubov coefficient. Another useful method is using the Virasoro anomaly.
III.1 Bogoliubov method
@@
The normalization constant $N_{\omega}$ is determined by requiring orthonormality of the modes under the Klein–Gordon inner product, defined on a constant-time hypersurface $\Sigma$ as [24, 6]
@@
for $l\neq m=1,2$. From this normalization condition, one obtains
@@
The right-moving component in wedge $R_{1}$ takes the form
@@
while in the remaining wedges, the expansions are
@@
III.1.1 Bogoliubov Transformation and the Number Operator
@@
III.2 Virasoro anomaly method
We consider a massless scalar field in two dimensions. We can therefore use the techniques of CFT in order to evaluate the particle content. We use the standard results involving Virasoro anomaly and the normal-ordered stress tensor [4]. In this work, we employ Rindler coordinates, which cover only a portion of Minkowski spacetime. Nonetheless, the normal-ordered stress tensor is well defined within the Rindler wedge and corresponds to renormalisation with respect to the Rindler vacuum [4, 3]. The resulting expectation values have interpretation in terms of observables accessible to uniformly accelerated observers and are directly related to the Unruh effect [20, 21, 22]. Under conformal transformation that maps from a coordinate chart $(U,V)$ to $(u,v)$ via the relation, $U=U(u)$ and similarly $V=V(v)$, the energy-momentum tensor behaves as
$T^{\prime}_{uu}=\bigg(\frac{\partial U}{\partial u}\bigg)^{2}T_{UU}+\frac{c}{12}S(U,u)$
Where $S(U,u)$ denotes the Schwarzian derivative and is defined as [3]. And, $c$ denotes the central charge. In the current context, the central charge is $c=-\hbar/2\pi$ [5].
@@ III.2 Virasoro anomaly method
So the relevant equation for the left-moving sector is,
$T^{\prime}_{uu}=\bigg(\frac{\partial U}{\partial u}\bigg)^{2}T_{UU}-\frac{\hbar}{24\pi}S(U,u)$
The corresponding equations for the $V$ sector read,
$T^{\prime}_{vv}=\bigg(\frac{\partial V}{\partial v}\bigg)^{2}T_{VV}-\frac{\hbar}{24\pi}S(V,v)$
where $S(V,v)$ is the Schwarzian derivative defined by replacing the $U,u$s with the $V,v$s.
@@ IV Particle content in $R_{3}$ along various paths
The explicit calculations and physical interpretations follow in subsequent subsections.
@@ IV.1 Particle content along path 1
In this subsection, we evaluate the particle content along Path 1, focusing on the particle content in $R_{3}$ when the field is in the Minkowski vacuum. The wedge $R_{3}$ is spatially shifted from the origin by a parameter $\Delta_{1}$, as illustrated in Figs. 1 and 2. Since the Unruh effect is invariant under such spatial translations, we expect the particle content in $R_{3}$ to be thermal. We demonstrate this explicitly below.
@@
Thus, the right-moving modes in wedge $R_{3}$, when evaluated in the Minkowski vacuum, are thermally excited and exhibit a Planck distribution. A similar calculation shows that the left-moving modes also thermalise, confirming full thermality along Path 1.
@@ IV.2 Particle content along path 2
In this section, we evaluate the particle content along Path 2, which connects the Rindler wedges $R_{1}$ and $R_{3}$ through a spatial shift parametrised by $\Delta_{1}$, as illustrated in Figs. 1 and 2. To compute the Bogoliubov coefficients relating the mode functions in these two wedges, we proceed as follows.
@@
To calculate the required Bogoliubov coefficients, we begin by multiplying both sides of Eq. (48) by the Fourier mode
$\int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}e^{i\Omega^{\prime}u_{2}}.$
After simplifications and using the orthogonality of exponentials, we obtain [14]
$\begin{split}\int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}\;e^{i\;\Omega^{\prime}\;u_{2}}\;\hat{\phi}_{2}&=\int_{0}^{\infty}\frac{d\Omega}{\sqrt{2\pi}}\frac{1}{\sqrt{2\Omega}}\\ \int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}\;\bigg(\overrightarrow{\hat{a}}_{2}(\Omega)\;e^{i\;(\Omega^{\prime}-\Omega)\;u_{2}}&+\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\;e^{i\;(\Omega+\Omega^{\prime})\;u_{2}}\bigg).\end{split}$ (95)
@@
due to orthogonality, valid for $\Omega^{\prime}>0$. Next, we compute the same integral using the field operator $\hat{\phi}_{1}(u_{1})$, expressed in terms of $u_{2}$ via the null-shift relation connecting wedges $R_{1}$ and $R_{2}$. Multiplying both sides of Eq. (47) by the same Fourier kernel gives
$\begin{split}\int_{-\infty}^{\infty}\frac{du_{2}}{\sqrt{2\pi}}\;e^{i\;\Omega^{\prime}\;u_{2}}\;\hat{\phi}_{1}=\frac{1}{2\pi}\int_{0}^{\infty}\frac{d\omega}{\sqrt{2\omega}}\bigg(\overrightarrow{\hat{a}}_{1}(\omega)\int_{-\infty}^{\infty}du_{2}\\ \;e^{i\;\Omega^{\prime}\;u_{2}-i\;\omega\;u_{1}}+\overrightarrow{\hat{a}}_{1}^{\dagger}(\omega)\;\int_{-\infty}^{\infty}du_{2}\;e^{i\;\Omega^{\prime}\;u_{2}+i\;\omega\;u_{1}}\bigg).\end{split}$ (97)
@@
The Rindler vacuum in wedge $\mathrm{R}_{1}$ is defined by the annihilation condition $\overleftarrow{\hat{b}}_{1}|0_{R_{1}}\rangle=0,$ and factorises as $|0_{R_{1}}\rangle=|\overleftarrow{0}_{R_{1}}\rangle\otimes|\overrightarrow{0}_{R_{1}}\rangle,$ where $|\overleftarrow{0}_{R_{1}}\rangle$ and $|\overrightarrow{0}_{R_{1}}\rangle$ denote the left- and right-moving vacuum states, respectively. The operators $\overleftarrow{\hat{b}}_{1}$ and $\overleftarrow{\hat{b}}_{1}^{\dagger}$ satisfy canonical commutation relations analogous to those in Eq. (60).
@@
This integral is analytically tractable (see Appendix. B) and evaluates to
@@
Here, $\delta(0)$ formally arises from the infinite volume limit and can be interpreted as a constant density over a finite spacetime volume [6]. Accordingly, the mean particle number density for frequency $\Omega$ is given by the Planckian distribution [11]:
@@
This confirms that the left-moving modes are thermally excited during the transition from $\mathrm{R}_{1}$ to $\mathrm{R}_{2}$, at a characteristic Unruh-like temperature $T=\frac{a}{2\pi}$. In contrast, the right-moving modes remain in the vacuum state. We therefore have the left-moving modes in a thermal state, and the right-moving modes have no excitations. This implies that the right-moving modes are in the ground state for all momenta $k$,
$|\overrightarrow{0_{k}\rangle}_{R_{1}}=|\overrightarrow{0_{k}\rangle}_{R_{2}}$ (113)
@@
Since $\mathrm{R}_{2}$ is obtained from $\mathrm{R}_{1}$ via a null shift, the resulting quantum field configuration exhibits selective thermalisation: the left-moving modes become thermally populated, while the right-moving modes remain in the vacuum, as previously established. We now compute the expectation value of the number operator for right-moving modes in $\mathrm{R}_{3}$ evaluated in the state $\mid\overrightarrow{0}_{R_{2}}\rangle$. The number operator
@@
$\langle\overrightarrow{\hat{N}}_{32}(\nu)\rangle=\langle\overrightarrow{0}_{R_{2}}\mid\overrightarrow{\hat{a}}_{3}^{\dagger}(\nu)\overrightarrow{\hat{a}}_{3}(\nu)\mid\overrightarrow{0}_{R_{2}}\rangle.$ (122)
@@
$\begin{split}\langle\overrightarrow{\hat{N}}_{32}(\nu)\rangle=\int_{0}^{\infty}d\Omega\int_{0}^{\infty}d\Omega^{\prime}\,\overrightarrow{\beta^{*}_{32}}(\nu,\Omega)\overrightarrow{\beta}_{32}(\nu,\Omega^{\prime})\\ \times\langle\overrightarrow{0}_{R_{2}}\mid\overrightarrow{\hat{a}}_{2}(\Omega^{\prime})\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\mid\overrightarrow{0}_{R_{2}}\rangle.\end{split}$ (123)
This implies,
$\begin{split}\langle\overrightarrow{0}_{R_{2}}\mid\overrightarrow{\hat{a}}_{2}(\Omega^{\prime})\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\mid\overrightarrow{0}_{R_{2}}\rangle=\langle\overrightarrow{0}\mid\overrightarrow{\hat{a}}_{2}(\Omega^{\prime})\overrightarrow{\hat{a}}_{2}^{\dagger}(\Omega)\mid\overrightarrow{0}\rangle\\ =\delta(\Omega-\Omega^{\prime}).\end{split}$ (124)
@@
The factor $\delta(0)$ arises from the mode normalisation in the continuum and represents an infinite constant related to the volume. Hence, the mean particle number density per unit frequency $\nu$ in $\mathrm{R}_{3}$, as measured in $\mathrm{R}_{2}$, is
@@
expressing the familiar Planckian distribution at the Unruh like temperature $T=\frac{a}{2\pi}$. We conclude that the right-moving modes mediating between $\mathrm{R}_{2}$ and $\mathrm{R}_{3}$ exhibit a thermal population characterised by this Planckian spectrum when the field in $\mathrm{R}_{2}$ is selectively excited.